Friday, 19 October 2012

Air is Down - Spectrometer out of service

The Air is down in TRACES lab.
Therefore, the Spectrometer will be out of service for the weekend until the problem is fixed.

Tuesday, 9 October 2012

Temperature Calibration

Dear User:

I am currently doing Temperature Calibration.
The spectrometer will be taken off automation for the time being.
I will let you all know when the calibration is finished.

N2 Fill Completed

Dear Users:

The N2 Fill is completed as schedule.  The spectrometer is ready for use.

Tuesday, 2 October 2012

NMR N2 filling - Done

Dear User:

The weekly N2 filling is done and the spectrometer is ready to use.